# Sampling 3D Molecular Conformers with Diffusion Transformers

### v1

@@ F.1 Index Positional Encoding (iPE)
Tab. A14 compares a variant including index positional encoding (iPE) from classic transformer architectures with our base model using aPE on QM9. Specifically, we use the node index and encode it via sinusoidal encodings into the tokens $\mathcal{H}$ before the first DiTMC block, similar to embedding the absolute positions via aPE. Since the graphs are generated from SMILES strings via rdkit and rdkit has to some extend a canonical ordering, this information can be used by the transformer architecture. However, index positional encoding breaks permutation equivariance (as we show in Tab. A14). This might be undesirable as permutation equivariance is one of the fundamental symmetries when learning on graphs. Since the ordering of atoms in a SMILES string is not uniquely defined, the trained network depends on the used framework for parsing the SMILES string or even a particular software version. We use rdkit (version 2024.9.5) for parsing SMILES strings to graphs.
@@
Nevertheless, our model using iPE can effectively exploit the information contained in atom indices assigned by rdkit. A version of our base aPE model without atom-pair conditioning but iPE achieves comparable performance to our model using geodesic distances as atom-pair conditioning (all-pair conditioning). As our all-pair conditioning strategy is similarly or more effective than iPE but additionally preserves permutation equivariance, it should be preferred over iPE and we don’t use iPE in any of our other experiments.
@@
Table A14: Ablating index positional encoding (iPE) on QM9 for different conditioning strategies (in brackets). To show the effect of atom permutations, we include results with randomly permuted atom indices (perm.). -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
@@
DiTMC+aPE-B (node & all-pair), perm. 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
F.2 Gaussian vs. Harmonic Prior
As shown in Tab. A15, using the harmonic prior improves all metrics slightly for our models on GEOM-QM9. Using the harmonic prior however doesn’t seem to be a crucial ingredient for the success of our method, as differences between Gaussian and Harmonic prior appear diminishing. As the results in Tab. A15 verify, our method can also be used with a simple Gaussian prior effectively. For larger molecular graphs the expensive eigendecomposition required for the Harmonic prior could therefore be avoided, which helps scaling our approach more easily.
Table A15: Ablation of PE strategies and Gaussian (G) and Harmonic (H) prior. We report mean coverage (COV) at a threshold of $0.5\mbox{\AA}$, and mean average minimum RMSD (AMR) for Recall -R and Precision -P. Best results in bold. All results are averaged over three random seeds.
PE strategy COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
G H G H G H G H
DiTMC+aPE 96.2 96.1 0.074 0.073 95.2 95.4 0.087 0.085
DiTMC+rPE 96.0 96.3 0.073 0.070 95.2 95.7 0.084 0.080
DiTMC+PE(3) 95.7 95.7 0.069 0.068 93.5 93.4 0.090 0.089
Appendix G Analysis of training loss as a function of latent time
In this section, we provide details for the analysis in Fig. 2C in the main part of the paper. We investigate the effect of the PEs and self-attention formulations on the accuracy of the model. We therefore take pre-trained models on GEOM-QM9 and compute the training loss (as detailed in Algorithm 1) averaged over 1000 samples drawn randomly from the GEOM-QM9 validation set. We compute the loss for 30 logarithmically spaced values of $t_{i}=1-10^{x_{i}}$, where $x_{i}\in[-1.8,0]$ with uniform spacing. We skip the stochastic term in the loss as is done while sampling from the ODE.
@@ Appendix G Analysis of training loss as a function of latent time
We further note, that absolute loss values for models trained with all our PE strategies decrease as latent time increases (see Fig. A4). This is expected, as conditional vector fields for each data sample will start to interact more strongly moving away from the data distribution. Our weighted loss (see Appendix A) effectively penalizes errors close to the data distribution during training and helps with keeping the error low in this important regime.
Figure A6: Loss as a function of time comparing different PE strategies. Results averaged over 1000 samples randomly drawn from the GEOM QM9 validation set. Left: loss relative to PE(3) as a baseline. In the important regime close to the data distribution, the model PE(3) has lower loss, yielding higher sample fidelity. Right: absolute loss values for all PEs. The loss decreases close to the data distribution for all models.
Appendix H SMILES Classification Experiment via Conditioning GNN
Table A16: We measure the discriminative power of our conditioning graph network on a training set of 1000 randomly sampled SMILES strings from the Geom QM9 validation set, as well as a toy dataset of 3 different SMILES strings. We investigate the required number of message passing layers, as well as using pre-trained weights from an end-to-end trained model.
GNN layers Weight init Trainable Accuracy (Geom QM9) Accuracy (Toy Data)
@@ Appendix H SMILES Classification Experiment via Conditioning GNN
A critical requirement of our DiTMC approach is the ability to disentangle distinct SMILES representations through conditioning. We investigate whether our proposed conditioning strategy is capable of learning the necessary information to distinguish between different SMILES strings for the generation of matching molecular conformers.
As a proxy evaluation task, we assess whether the conditioning network alone can function as a classifier of SMILES strings. To this end, we construct two training datasets: a toy dataset comprising three specific SMILES strings of a hydroxyl group moving along a carbon chain (C(O)CCCCCCCC, CC(O)CCCCCCC, CCC(O)CCCCCC), and a larger set consisting of 1000 randomly sampled SMILES strings drawn from the GEOM-QM9 validation set. Each SMILES string becomes a seperate class, so for each class there is exactly one example in the training data. The classification task is performed on the graph representations of the molecules, employing the same feature set and GNN architecture utilized in our conditioning graph neural network (Sec. 4.1, appendix D.2) plus a simple classification head.
We train models with a batch size of 3 for 5000 epochs (toy dataset) and batch size of 64 over 250 epochs (GEOM-QM9 subset). We report classification accuracy on the training sets directly to evaluate the model’s discriminative capacity. Furthermore, we explore whether conditioning weights obtained from an end-to-end trained model retain discriminatory power by freezing them and attaching a linear classification head.
Appendix I Architecture Details for Conditioning GNN
To transform SMILES representations to per-token conditioning information, we use a GNN that directly operates on the bond graph induced by the SMILES string. Our model processes molecular graphs by first embedding node and edge features, initially represented as one-hot vectors (a full list of features is provided in Tab. A7). These features are projected into a shared latent space using two-layer multilayer perceptrons (MLPs). For message passing, we employ a graph neural network architecture inspired by the processor described in the MeshGraphNet (MGN) framework [54].
The GNN maintains and updates both node and edge representations across multiple layers. Each message passing block consists of two main steps: first, the edge representations are updated based on the current edge representations and the representations of the connected nodes:
The final output of the described GNN is a set of node embeddings per atom, and a set of edge embeddings per bond. Those serve as atom-wise or bond-pair conditioning inputs to the DiT transformer, as detailed in Sec. 4.1.
Appendix J Autoguidance
We apply autoguidance, a recent technique for enhancing generative models by leveraging the outputs of a weaker model to guide sampling from a stronger one [71]. Originally proposed in the context of image diffusion, autoguidance works by introducing a degraded model with parameters $\hat{\theta}$ trained on the same data and conditioning as the main model with parameters $\theta$, but with additional constraints such as reduced capacity, increased noise or fewer gradient steps. During generation, the predicted vector field $\boldsymbol{v}^{\theta}(\boldsymbol{x},t)$ is guided using the vector field predicted by $\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where the discrepancy between the two models serves as a corrective signal. Intuitively, this nudges the generation process towards higher-quality samples by amplifying the difference between a less accurate and a more accurate model.
In our molecular conformer generation setup, we choose the parameters after training for 5 epochs without applying exponential moving average (EMA) for $\hat{\theta}$. We then use the difference in predicted vector fields from to bias the conformer generation trajectory. Following Ref. [72], we apply autoguidance only in the guidance interval $t\in[0,t_{max}]$ and additionally ablate $t_{max}$ as the upper bound of the guidance interval. It becomes straightforward to adapt Algorithm 2 to reflect these changes during sampling as a drop-in repacement for the predicted vector field. The guided velocity (for $t\in[0,t_{max}]$) is computed as $\boldsymbol{v}_{\eta}(\boldsymbol{x},t)=\eta\boldsymbol{v}^{\theta}(\boldsymbol{x},t)+(1-\eta)\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where $\eta$ is the guidance strength.
Fig. A7 shows that autoguidance yields enhanced fidelity of generated conformations, as indicated by improvements in precision metrics, but might negatively impact diversity (recall). Increasing guidance strength improves precision (COV-P, AMR-P) but may reduce recall (COV-R, AMR-R). A similar trend can be observed for the guidance interval, where increased $t_{max}$ leads to stronger effects of autoguidance. This effectively establishes a trade-off between precision and recall that can be tuned post-hoc after training with minimal effort.
Figure A7: Autoguidance performance on GEOM-QM9 using our DiTMC+aPE-B model. We evaluate the effect of autoguidance on conformer generation quality across various guidance intervals ($t_{max}$) on the x-axes, and guidance strength ($\eta$) as different colors. Dashed black lines indicate performance without guidance.
Appendix K Analysis of Sampling Trajectories
@@ Appendix I Architecture Details for Conditioning GNN
Figure A8: Comparison of molecular generation with node-only versus node- and all-pair conditioning on GEOM-QM9. Each row shows a prior sample, sampling trajectories, and final generated structure for both models; pairwise conditioning preserves bonds from the conditioning graph $\mathcal{G}$.
@@
Fig. A8 compares the generative performance of DiTMC models trained on the GEOM-QM9 dataset under two different conditioning strategies: (1) node-only conditioning and (2) node plus pairwise conditioning. Each row in the figure corresponds to one example molecule selected from the test set. The molecules are chosen to maximize the root mean squared deviation (RMSD) between the final generated structures of the two models. This selection highlights cases where the differences between the conditioning schemes are most pronounced.
@@
Within each row, we display a sequence of images: the initial prior sample, followed by the intermediate trajectory of the ODE sampling process over 50 sampling steps for the node-only conditioned model, and the resulting final structure (with predicted bonds rendered in yellow). This sequence is repeated for the node- and all-pair-conditioned model, allowing a side-by-side visual comparison of the generation dynamics and final outputs.
@@
The results reveal a consistent pattern: models trained with node-only conditioning fail to preserve bonding patterns from the conditioning graph $\mathcal{G}$. This manifests as bond stretching or atom permutation in the final structure. In contrast, the pairwise-conditioned model produces geometries that adhere more closely to expected chemical structure and the given bonds. We note that if atoms are simply permuted by the model using only node conditioning, the generated structure might still be valid in terms of the combination of generated 3D positions and atom types. The degraded performance of node conditioning versus node and pair-conditioning can therefore in part be explained by the used RMSD and Coverage metrics, which are not invariant to permutations of atoms.

### v2

@@ I.3 Index Positional Encoding (iPE)
Tab. A15 compares a variant including index positional encoding (iPE) from classic transformer architectures with DiTMC+aPE-B on GEOM-QM9. Specifically, we use the node index and encode it via sinusoidal encodings into the tokens $\mathcal{H}$ before the first DiTMC block, similar to embedding the absolute positions via aPE. Since the molecular graphs are generated from SMILES strings via RDKit and RDKit has to some extend a canonical ordering, this information can be used by the transformer architecture. However, index positional encoding breaks permutation equivariance (as we show in Tab. A15). This might be undesirable as permutation equivariance is one of the fundamental symmetries when learning on graphs. Since the ordering of atoms in a SMILES string is not uniquely defined, the trained network depends on the used framework for parsing the SMILES string or even a particular software version. We use RDKit (version 2024.9.5) for parsing SMILES strings to graphs.
@@
Nevertheless, our DiTMC+aPE-B using iPE can effectively exploit the information contained in atom indices assigned by RDKit. A version of DiTMC+aPE-B without atom-pair conditioning but iPE achieves comparable performance to DiTMC+aPE-B using geodesic distances as atom-pair conditioning (pairwise conditioning). As our pairwise conditioning strategy is similarly or more effective than iPE but additionally preserves permutation equivariance, it should be preferred over iPE and we don’t use iPE in any of our other experiments.
@@
Table A15: Ablating index positional encoding (iPE) on GEOM-QM9 for different conditioning strategies (in brackets). To show the effect of atom permutations, we include results with randomly permuted atom indices (perm.). -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
@@
DiTMC+aPE-B (node & pairwise) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
DiTMC+aPE-B (node & pairwise), perm. 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
Appendix J Analysis of training loss as a function of latent time
In this section, we provide details for the analysis in Fig. 2C in the main part of the paper. We investigate the effect of the positional embeddings and self-attention formulations on the accuracy of the model. We therefore take pre-trained models on GEOM-QM9 and compute the training loss (as detailed in Algorithm 1) averaged over 1000 samples drawn randomly from the GEOM-QM9 validation set. We compute the loss for 30 logarithmically spaced values of $t_{i}=1-10^{x_{i}}$, where $x_{i}\in[-1.8,0]$ with uniform spacing. We skip the stochastic term in the loss as is done while sampling from the ODE.
@@
We further note, that absolute loss values for models trained with all our PE strategies decrease as latent time increases (see Fig. A5). This is expected, as conditional vector fields for each data sample will start to interact more strongly moving away from the data distribution. Our weighted loss (see Appendix B) effectively penalizes errors close to the data distribution during training and helps with keeping the error low in this important regime.
Figure A5: Loss as a function of time comparing different PE strategies. Results averaged over 1000 samples randomly drawn from the GEOM-QM9 validation set. Left: loss relative to PE(3) as a baseline. In the important regime close to the data distribution, the model PE(3) has lower loss, yielding higher sample fidelity. Right: absolute loss values for all PEs. The loss decreases close to the data distribution for all models.
Appendix K SMILES Classification Experiment via Conditioning GNN
The molecular graphs in our dataset are represented as SMILES strings. A critical requirement of DiTMC is the ability to distinguish distinct molecular graphs or SMILES representations through conditioning. In this section, we investigate whether our conditioning GNN is capable of learning the necessary information to distinguish between different SMILES strings for the generation of matching molecular conformers.
As a proxy evaluation task, we assess whether the conditioning network alone can function as a classifier of SMILES strings. To this end, we construct two training datasets: a toy dataset comprising three specific SMILES strings of a hydroxyl group moving along a carbon chain (C(O)CCCCCCCC, CC(O)CCCCCCC, CCC(O)CCCCCC), and a larger set consisting of 1000 randomly sampled SMILES strings drawn from the GEOM-QM9 validation set. Each SMILES string becomes a seperate class, so for each class there is exactly one example in the training data. The classification task is performed on the graph representations of the molecules, employing the same feature set and GNN architecture utilized in our conditioning GNN (see Appendix D.3 and Appendix G.2) followed by a simple classification head.
We train different models with a batch size of 3 for 5000 epochs on the curated toy dataset and batch size of 64 over 250 epochs on the GEOM-QM9 subset. We report classification accuracy on the training sets directly to evaluate the model’s discriminative capacity. Furthermore, we explore whether conditioning weights obtained from an end-to-end trained model retain discriminatory power by freezing them and attaching a linear classification head.
Table A16: We measure the discriminative power of our conditioning graph network on a training set of 1000 randomly sampled SMILES strings from the GEOM-QM9 validation set, as well as a toy dataset of 3 different SMILES strings. We investigate the required number of message passing layers, as well as using pre-trained weights from an end-to-end trained model.
GNN layers Weight init Trainable Accuracy (GEOM-QM9) Accuracy (Toy Data)
@@
Appendix L Analysis of Sampling Trajectories
@@
Figure A7: Comparison of molecular generation with node-only versus node- and pair-wise (all-pair) conditioning on GEOM-QM9. Each row shows a prior sample, sampling trajectories, and final generated structure for both models; pairwise conditioning preserves bonds from the conditioning graph $\mathcal{G}$.
@@
Fig. A7 compares the generative performance of DiTMC models trained on the GEOM-QM9 dataset under two different conditioning strategies: (1) node-only conditioning and (2) node plus pairwise conditioning, where we use the all-pair conditioning based on geodesic graph distances. Each row in the figure corresponds to one example molecule selected from the test set. The molecules are chosen to maximize the root mean squared deviation (RMSD) between the final generated structures of the two models. This selection highlights cases where the differences between the conditioning schemes are most pronounced.
@@
Within each row, we display a sequence of images: the initial prior sample, followed by the intermediate trajectory of the ODE sampling process over 50 sampling steps for the node-only conditioned model, and the resulting final structure (with predicted bonds rendered in yellow). This sequence is repeated for the node- and pair-wise conditioned model, allowing a side-by-side visual comparison of the generation dynamics and final outputs.
@@
The results reveal a consistent pattern: models trained with node-only conditioning fail to preserve bonding patterns from the conditioning graph $\mathcal{G}$. This manifests as bond stretching or atom permutation in the final structure. In contrast, the model, that is conditioned on pairwise geodesic distances, produces geometries that adhere more closely to expected chemical structure and the given bonds. We note that if atoms are simply permuted by the model using only node conditioning, the generated structure might still be valid in terms of the combination of generated 3D positions and atom types. The degraded performance of node conditioning versus node and pair-conditioning can therefore in part be explained by the used RMSD and Coverage metrics, which are not invariant to permutations of atoms.
@@
Appendix N Code and Data Availability
The code and data to reproduce the main results of this paper, can be downloaded from here: https://doi.org/10.5281/zenodo.15489212, or via github: https://github.com/ML4MolSim/dit_mc.
Figure A8: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+aPE and other state-of-the-art methods on GEOM-DRUGS. The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
Figure A9: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+rPE and other state-of-the-art methods on GEOM-DRUGS. The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
Figure A10: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+PE(3) and other state-of-the-art methods on GEOM-DRUGS. The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
Figure A11: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer. For each model, markers from left to right correspond to an increasing number of sampling steps during generation. Here, we follow Refs. [hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.
Figure A12: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer. For each model, markers from left to right correspond to an increasing number of sampling steps during generation. Here, we follow Refs. [hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.
Figure A13: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer. For each model, markers from left to right correspond to an increasing number of sampling steps during generation. Here, we follow Refs. [hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.